Compare the calculated zero-point energies according to Quantum Mechanics and Classical Mechanics for a xenon atom confined to a 10 nm line.
Answer #1

according to Quantum Mechanics

ZPE = h2/8ml2 = (6.626 x 10-34)2 /{8 x 9.1 x 10-31 x (10 x 10-9)2} =  6.0308 x 10-22 J  

According to according to Quantum Mechanics ZPE = 0 (energy at 0 K is ZPE)